A student is asked to standardize a solution of barium hydroxide. He weighs out 1.07 g potassium hydrogen phthalate (KHC8H4O4, treat this as a monoprotic acid). It requires 29.8 mL of barium hydroxide to reach the endpoint. A. What is the molarity of the barium hydroxide solution? M This barium hydroxide solution is then used to titrate an unknown solution of hydrochloric acid. B. If 10.4 mL of the barium hydroxide solution is required to neutralize 10.8 mL of hydrochloric acid, what is the molarity of the hydrochloric acid solution?
Answer #1

C-OH 920 O=0 ure + Ba(OH), Ba2+ + 2420 4-0- v=0 Polanium hydrogen phthalate Given weight 1 киP = (-ona molar manj kHP = 204.Molarity 9 BacOH)2 = 0.0879 M B. Bacol)a +246 -→ mol of BacoH)ą used = BaCl2 + 2420 molarity & volume = 0.0879 mol x 0.0104x